A sum of $2100 is invested in two accounts, one earning 7% simple annual interest and one earning 4%. If the total interest earned for the year is $120, how much money is in each account?
Answer by math problem solving(30)

Let x be the amount of money invested at 7%.
Then the amount invested at 4% is 2100 - x.
The interest earned from the money invested at 7% is x*7%
The interest earned from the money invested at 4% is (2100 - x)*4%
As the total interest earned from the two investments is given as $120, we have:
x*7% + (2100 - x)*4% = 120
Solving for x, we get
7x + 4(2100 - x) = 12000 (multiply both sides by 100)
7x + 8400 - 4x = 12000
3x + 8400 = 12000
3x = 3600
x = 1200
So $1200 is invested at 7%, and $2100 - $1200 = $900 is invested at 4%.